OpenStudy (anonymous):

Can someone show how how to properly solve this problem, the directions are to Find the derivative of the function: s(x)= ln(4x-2)^1.1

OpenStudy (anonymous):

I know that the derivative of (4x-2)^1.1 equals 1.1*(4)^0.1, but the "ln" is what's confusing me.

OpenStudy (kropot72):

Let (4x - 2)^1.1 = z Then s(x) = ln z \[\frac{ds}{dz}=\frac{1}{z}\] \[\frac{ds}{dx}=\frac{ds}{dz}.\frac{dz}{dx}=\frac{1}{z}.\frac{d(4x-2)^{1.1}}{dx}\]

OpenStudy (kropot72):

\[\frac{ds}{dx}=\frac{4.4(4x-2)^{0.1}}{(4x-2)^{1.1}}=\frac{4.4}{4x-2}\]

OpenStudy (anonymous):

Oh! Okay, I got it. Thank you soo much =)

OpenStudy (kropot72):

You're welcome :)